Tell me how to solve a problem on geometry .there is a question for finding the radius of circle from the given equation
equation like : x^2+4x-y^2+5y=12

Grade:Upto college level

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

equation of circle

x 2 + y 2 + a x + b y = c

has center has coordinates (-a/2 , -b/2) and radius r such that r 2 = c + (a/2) 2 + (b/2) 2

substitute values to get the answer !!